OpenStudy (anonymous):

Find the area of the region bound on the left by the line x = 1, on the right by x = 2, the x-axis, and the curve.

OpenStudy (anonymous):

\[\frac{ x-3 }{ x^3+x^2}\]

OpenStudy (anonymous):

is the equation of the curve

OpenStudy (anonymous):

\[\int\limits_{1}^{2}(-\frac{ x-3 }{ x^3+x^2 })dx\]

OpenStudy (anonymous):

Is this right?

OpenStudy (anonymous):

Where did the negative sign come from?

OpenStudy (anonymous):

0-the equation because the x axis is on top

OpenStudy (anonymous):

Oh, right! This integral requires partial fraction decomposition. \[-\frac{x-3}{x^3+x^2}=-\frac{x-3}{x^2(x+1)}=-\frac{A}{x}-\frac{B}{x^2}-\frac{C}{x+1}\] \[x-3=Ax(x+1)+B(x+1)+Cx^2\\ x-3=(A+C)x^2+(A+B)x+B\] Matching up coefficients, you have \[\begin{cases}A+C=0\\A+B=1\\B=-3\end{cases}\] Solve for the constants, plug them into the above fractions, then integrate term by term.

OpenStudy (anonymous):

so B=-3, A=4, C=-4

OpenStudy (anonymous):

\[-(4lnx-3lnx^2-4\ln(x+1))\]

OpenStudy (anonymous):

right?

OpenStudy (alekos):

2nd term should be -3x^-1

OpenStudy (anonymous):

oh right! thank you

OpenStudy (alekos):

i'm not sure about the -ve signs for each term. doesnt look right

OpenStudy (anonymous):

I got the right answer

OpenStudy (alekos):

ok good. so the signs are correct, well done